OpenStudy (anonymous):

During a titration, the following data were collected. A 25 mL portion of an unknown monoprotic acid solution was titrated with 1.0 M NaOH(aq). 65 mL of base solution were required to neutralize the sample. How many moles of acid are present in 3.0 L of this unknown solution?

OpenStudy (aaronq):

NaOH and the monoprotic acid are in a 1:1 stoichiometric ratio in the neutralization (acid-base) reaction, meaning that 1 mole of base neutralizes 1 mole of acid. use: \(M_{acid}V_{acid}=M_{base}V_{base}\), to find the molarity of the acid. then use: \(M_{acid}=\dfrac{n_{acid}}{L_{solution}}\), to find the moles (n) of acid.

OpenStudy (anonymous):

its a multiple choice question and I can't seem to be getting the answer. The choices are -2.6 moles -0.9 moles -8.7 moles -7.8 moles -5.6 moles

OpenStudy (aaronq):

(0.025 L *\(M_{acid}\)=1.0 M NaOH*0.065 L) \(M_{acid}\)=\(\dfrac{n_{acid}}{L_{solution}}=2.6 M=\dfrac{n_{acid}}{3\;L}\rightarrow3L*2.6M=7.8\;moles\)